Lou Barlow, a divisional manager for Sage Company, has an opportunity to manufacture and sell one of two new products for a five-year period. His annual pay raises are determined by his division’s return on investment (ROI), which has exceeded 19% each of the last three years. He has computed the cost and revenue estimates for each product as follows:

Product A Product B
Initial investment:
Cost of equipment (zero salvage value) $ 190,000 $ 400,000
Annual revenues and costs:
Sales revenues $ 270,000 $ 370,000
Variable expenses $ 128,000 $ 178,000
Depreciation expense $ 38,000 $ 80,000
Fixed out-of-pocket operating costs $ 72,000 $ 52,000

The company’s discount rate is 17%.

Required:

1. Calculate the payback period for each product.

2. Calculate the net present value for each product.

3. Calculate the internal rate of return for each product.

4. Calculate the project profitability index for each product.

5. Calculate the simple rate of return for each product.

6a. For each measure, identify whether Product A or Product B is preferred.

6b. Based on the simple rate of return, Lou Barlow would likely:
Answer #1

Computation of Annual cash Inflow:

Product A Product B
Sales revenue 270,000 370,000
Less:
Variable expenses (128,000) (178,000)
Fixed out of pocket operating cost (72,000) (52,000)
Annual cash inflows 70,000 140,000

Note: As tax are not given here, so depreciation will not be used for calculation of annual cash flows.

1 1. Payback Period 2 Initial Investment 3 Annual Cash inflows 4 Payback Period Product A 190000 70000 2.71 Product B 400000

28 5. Annual Operating Income 29 Annual Cash inflows 30 Less: Depreciation expenses 70000 38000 32000 140000 80000 60000 31 3

Calculated as:

1 1. Payback Period 2 Initial Investment 3 Annual Cash inflows 4 Payback Period Product A 190000 =270000-128000-72000 =C2/C328 5. Annual Operating Income 29 Annual Cash inflows 30 Less: Depreciation expenses 70000 38000 =B29-B30 140000 80000 =C29-C3

6a.

Solution 6a:

Product Preference
Payback Period Product A
Net Present Value Product B
IRR Product A
Profitability index Product A
Simple rate of return Product A

Solution 6b:

Based on simple rate of return, Lou Barlow would likely to reject both the products as its less than 19%.